lizzie777:

Complete the equation so it has infinitely many solutions. 4x + 7 = 4(x + 3) –

lizzie777:

help

justus:

To have an infinite number of solutions in this equation, we must solve for x and reach a point where we have "0x = 0."

lizzie777:

ok

lizzie777:

what do we do first

justus:

The missing term which we will call it 'y' is in the equation 4x + 7 = 4(x + 3) – y

justus:

Next you have to distribute 4 over the parentheses on right side 4x + 7 = 4x + 12 – y

justus:

Then substract 7 on both sides 4x + 7 - 7 = 4x + 12 – y - 7 4x = 4x + 5 – y

justus:

You now have minus 4x on both sides

justus:

4x - 4x = 4x + 5 – y - 4x

justus:

tell me what you get now

justus:

cancel out all the 4x terms and you will get it

justus:

@lizzie777 Are you there?

lizzie777:

yes

lizzie777:

is it 5

justus:

No, cancel out all the 4x terms in this equation and tell me what you are left with 4x - 4x = 4x + 5 – y - 4x

lizzie777:

5 – y

justus:

Correct

lizzie777:

so the missing piece is 0

lizzie777:

wait no sorry

lizzie777:

5 I mean

justus:

0x = 5 - y the equation will have infinitely many solutions if the right side is 0 5 - y = 0 5 = y

lizzie777:

the answer is 5

justus:

Yes.

lizzie777:

thanks